OpenStudy (scarlettfarra2000):

What am I doing wrong? Preform the indicated operation. Assume all variables represent nonnegative real numbers

OpenStudy (scarlettfarra2000):

\[\sqrt{6x ^{2}}+x \sqrt{52}\]

OpenStudy (scarlettfarra2000):

use the product rule to simplify the first radical \[\sqrt{6x ^{2}}=\sqrt{x ^{2}}\times \sqrt{6}=x \sqrt{6}\]

OpenStudy (scarlettfarra2000):

For the second radical factor 54 \[x \sqrt{54}=x \sqrt{9\times6}\]

OpenStudy (scarlettfarra2000):

use the product rule \[x \sqrt{9\times6}=x \sqrt{9}\times \sqrt{6}\] \[x \sqrt{9\times6}=3x \sqrt{9}\]

hartnn (hartnn):

you mean \(3x \sqrt 6\) right?

OpenStudy (scarlettfarra2000):

combine both radicals \[(9x+3)\sqrt{9}=3x \sqrt{9}\]

hartnn (hartnn):

also 6*9 is not 52

hartnn (hartnn):

oh just a typo

OpenStudy (scarlettfarra2000):

It's 54 I messed up

hartnn (hartnn):

\(x\sqrt 6 + 3x\sqrt 6 = (x+3x)\sqrt 6 = 4x \sqrt 6 \)

hartnn (hartnn):

\(\color{blue}{\text{Originally Posted by}}\) @ScarlettFarra2000 use the product rule \[x \sqrt{9\times6}=x \sqrt{9}\times \sqrt{6}\] \(x \sqrt{9\times6}=3x \sqrt{9}\) <<<<<< \(\color{blue}{\text{End of Quote}}\) \(x \sqrt{9\times 6} = x \sqrt 9 \sqrt 6 = 3x\sqrt 6 \)

OpenStudy (scarlettfarra2000):

Oh so its not \[3x \sqrt{9}\] like i thought its \[3x \sqrt{6}\]

hartnn (hartnn):

\(\Large x \sqrt{9\times 6} = x \sqrt 9 \sqrt 6 = 3x\sqrt 6 \) ^^

hartnn (hartnn):

doubt cleared? any more doubts?

OpenStudy (scarlettfarra2000):

Well when I put the answer in on my assignment to still tells me I'm wrong

hartnn (hartnn):

\(4x \sqrt 6\) is correct...

OpenStudy (scarlettfarra2000):

Yes thank you so much

hartnn (hartnn):

welcome ^_^